# Koby Hayashi

### v1

@@ Abstract
Symmetric Nonnegative Matrix Factorization (SymNMF) is a technique in data analysis and machine learning that approximates a symmetric matrix with a product of a nonnegative, low-rank matrix and its transpose. To design faster and more scalable algorithms for SymNMF we develop two randomized algorithms for its computation. The first algorithm uses randomized matrix sketching to compute an initial low-rank input matrix and proceeds to use this input to rapidly compute a SymNMF. The second algorithm uses randomized leverage score sampling to approximately solve constrained least squares problems. Many successful methods for SymNMF rely on (approximately) solving sequences of constrained least squares problems. We prove theoretically that leverage score sampling can approximately solve nonnegative least squares problems to a chosen accuracy with high probability. Finally we demonstrate that both methods work well in practice by applying them to graph clustering tasks on large real world data sets. These experiments show that our methods approximately maintain solution quality and achieve significant speed ups for both large dense and large sparse problems.
@@ 1 Introduction
We propose the first randomized algorithms for Symmetric Nonnegative Matrix Factorization (SymNMF). Nonnegative Matrix Factorization (NMF) is an important method in data analysis with applications to data visualization, text mining, feature learning, information fusion and more [28, 25, 46, 22, 11]. SymNMF is a variant of NMF where the input matrix is symmetric and the output low-rank approximation is also constrained to be symmetric [25, 49]. Applications of SymNMF include (hyper)graph clustering, image segmentation, and information fusion [44, 10, 19, 5, 6]. Several randomized algorithms for nonsymmetric NMF have been previously proposed and shown to be effective for dense and small sparse problems [43, 41, 13], but as far as we are aware there is no prior work on randomized algorithms for SymNMF. Our contributions in this work include a randomized algorithm for SymNMF we call “Low-rank Approximated Input SymNMF” (LAI-SymNMF), a randomized algorithm based on leverage score sampling for least squares problems we call LvS-SymNMF, novel theoretical analysis of leverage score sampling for the Nonnegative Least Squares problem and theoretical analysis of a hybrid sampling scheme for leverage score sampling.
@@
The rest of the paper is organized as follows. Section 2, which discusses background material including non-randomized SymNMF algorithms, reviews existing randomized NMF methods and other related work such as randomized methods for other low-rank matrix decompositions and tensor decompositions. Section 3 introduces our first proposed algorithm LAI-SymNMF, where LAI stands for Low-rank Approximate Input. LAI-SymNMF uses randomized techniques to rapidly compute an initial, unconstrained low-rank approximation before proceeding to compute an NMF of this LAI. Section 4 presents a row-sampling algorithm called LvS-SymNMF. This method solves a sequence of nonnegative least squares (NLS) problems using a technique called leverage score sampling to accelerate the solver. We use a hybrid approach, involving both deterministic and randomized sampling based on the leverage scores. Our theoretical analysis of this strategy gives the sample complexity needed to achieve an accuracy guarantee relative to the NLS residual, with high probability, and we empirically show its advantage over purely randomized sampling. Section 5 presents experimental results for the proposed algorithms on two real world data sets. Each data set is presented as a graph and the SymNMF output is used to cluster the vertices. Both methods achieve significant speed ups over deterministic methods ranging from $5$ to 7.5×7.5$\times$7.5 × and are able to maintain accuracy in terms of normalized residual norms and cluster quality.
2 Preliminaries and Related Work
@@
We begin by briefly discussing the NMF and SymNMF problems followed by an introduction to various methods for computing NMF’s and SymNMF’s. Then we discuss the “Randomized Range Finder” and leverage score sampling for least squares problems. Various notation is given in Table 1.
@@
${\bm{\mathbf{{I}}}}$ Identity Matrix ${\bm{\mathbf{{a}}}}_{i}$ $i$’th column vector of ${\bm{\mathbf{{A}}}}$
@@
${\bm{\mathbf{{A}}}}^{\sf T}$ Transposition $\mathbb{R}_{\geq 0}$ Nonnegative Reals
@@
$\min_{\{{\bm{\mathbf{{W}}}},{\bm{\mathbf{{H}}}}\}\geq 0}\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}$ (1)
where ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{m\times n}$, ${\bm{\mathbf{{W}}}}\in\mathbb{R}_{\geq 0}^{m\times k}$, and ${\bm{\mathbf{{H}}}}\in\mathbb{R}_{\geq 0}^{n\times k}$. The notation $\mathbb{R}_{\geq 0}$ denotes the nonnegative orthant and $k$ is the desired reduced rank, given as an input, and usually $k\ll\min(m,n)$. While ${\bm{\mathbf{{X}}}}$ is often also nonnegative, it is not strictly required. Since NMF is a non-convex optimization problem, optimization algorithms often find a local minimum at best. The works [16, 23] give a comprehensive discussion of NMF.
When ${\bm{\mathbf{{X}}}}$ is symmetric, ${\bm{\mathbf{{X}}}}={\bm{\mathbf{{X}}}}^{\sf T}$, it is often desired the two low-rank factors in Equation 1 be the same [25, 42, 49]. This problem is called Symmetric NMF (SymNMF) and its objective function is expressed as
$\min_{{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{H}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}.$ (2)
@@ 2.1 Algorithms for SymNMF
Many algorithms for solving Equation 2 have been proposed. Broadly these methods can be put into two categories: Alternating Updating (AU) methods and all-at-once optimization methods. AU methods alternate between updating a subset of the variables while holding others fixed, eventually iterating through the entire subset of variables. All-at-once methods update all of the variables simultaneously.
AU methods for SymNMF include symmetrically regularized Alternating Nonnegative Least Squares (ANLS) [25], symmetrically regularized Hierarchical Least Squares (HALS) [49], Cyclic Coordinate Descent (CCD) [42], and Progressive Hierarchical Alternating Least Squares (PHALS) [20]. All-at-once methods include Projected Gradient Descent (PGD), Projected Newton-like update [25], and Projected Gauss-Newton with Conjugate Gradients (PGNCG) [15]. This work focuses on methods based on regularized ANLS and the PGNCG algorithm. This is because the ANLS method is generally superior to both PGD and the Newton-like method [25], and the CCD method tends to be unsuitable for large data sets as it sequentially iterates over the elements of ${\bm{\mathbf{{H}}}}$. We now present the methods based on regularized ANLS and the PGNCG algorithm.
@@ 2.1.1 Symmetrically Regularized Alternating Nonnegative Least
$\min_{\{{\bm{\mathbf{{W}}}},{\bm{\mathbf{{H}}}}\}\geq 0}\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}^{2}+\alpha\|{\bm{\mathbf{{W}}}}-{\bm{\mathbf{{H}}}}\|_{F}^{2},$ (3)
where ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ are forced to be close to each other in Frobenius norm using a large value of $\alpha$, thus guiding the iterates towards a symmetric approximation. The authors of [49] show that under mild assumptions the critical points of Equation 2 and Equation 3 are the same and argue that this makes Equation 3 an appropriate surrogate for Equation 2. Equation 3 can be iteratively approximated by using the ANLS method for updating ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ as in the following two equations :
$\min_{{\bm{\mathbf{{W}}}}\geq 0}\Big{\|}\begin{bmatrix}{\bm{\mathbf{{H}}}}\\ \sqrt{\alpha}{\bm{\mathbf{{I}}}}\end{bmatrix}{\bm{\mathbf{{W}}}}^{\sf T}-\begin{bmatrix}{\bm{\mathbf{{X}}}}\\ \sqrt{\alpha}{\bm{\mathbf{{H}}}}^{\sf T}\end{bmatrix}\Big{\|}_{F}\ \text{and}\ \min_{{\bm{\mathbf{{H}}}}\geq 0}\Big{\|}\begin{bmatrix}{\bm{\mathbf{{W}}}}\\ \sqrt{\alpha}{\bm{\mathbf{{I}}}}\end{bmatrix}{\bm{\mathbf{{H}}}}^{\sf T}-\begin{bmatrix}{\bm{\mathbf{{X}}}}\\ \sqrt{\alpha}{\bm{\mathbf{{W}}}}^{\sf T}\end{bmatrix}\Big{\|}_{F}.$ (12)
@@
The ANLS method is generally superior to PGD and the Newton-like method [25]. This is mainly because PGD suffers from slow convergence and the Newton-like method is expensive for even moderately sized problems due to the need for approximate Hessian inversion.
@@ 2.1.2 Hierarchical Alternating Least Squares for SymNMF
A method that uses the HALS framework for optimizing Equation 3 was proposed in [49]. The update rules are given by the following two equations
${\bm{\mathbf{{w}}}}_{i}=\max\Big{(}\frac{({\bm{\mathbf{{R}}}}_{i}+\alpha{\bm{\mathbf{{I}}}}){\bm{\mathbf{{h}}}}_{i}}{\|{\bm{\mathbf{{h}}}}_{i}\|_{2}^{2}+\alpha},0\Big{)}\ \text{and}\ {\bm{\mathbf{{h}}}}_{i}=\max\Big{(}\frac{({\bm{\mathbf{{R}}}}_{i}+\alpha{\bm{\mathbf{{I}}}}){\bm{\mathbf{{w}}}}_{i}}{\|{\bm{\mathbf{{w}}}}_{i}\|_{2}^{2}+\alpha},0\Big{)}.$ (13)
@@
Here ${\bm{\mathbf{{W}}}}=\big{[}{\bm{\mathbf{{w}}}}_{1},\cdots,{\bm{\mathbf{{w}}}}_{k}\big{]}$, ${\bm{\mathbf{{H}}}}=\big{[}{\bm{\mathbf{{h}}}}_{1},\cdots,{\bm{\mathbf{{h}}}}_{k}\big{]}$ are the columns of ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ and ${\bm{\mathbf{{R}}}}_{i}={\bm{\mathbf{{X}}}}-\sum_{j\neq i,j=1}^{k}{\bm{\mathbf{{w}}}}_{j}{\bm{\mathbf{{h}}}}_{j}^{\sf T}.$ Following these rules, the columns of ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ are updated as pairs in sequence as ${\bm{\mathbf{{w}}}}_{1},{\bm{\mathbf{{h}}}}_{1},\cdots{\bm{\mathbf{{w}}}}_{i},{\bm{\mathbf{{h}}}}_{i},\cdots{\bm{\mathbf{{w}}}}_{k},{\bm{\mathbf{{h}}}}_{k}$. This method is less efficient as it relies on the matrix ${\bm{\mathbf{{R}}}}_{i}$, as in Equation 13.
To make the updates efficient we modify the update rule to
${\bm{\mathbf{{w}}}}_{i}\leftarrow\Big{(}\frac{({\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}+\alpha{\bm{\mathbf{{I}}}}){\bm{\mathbf{{h}}}}_{i}}{\|{\bm{\mathbf{{h}}}}_{i}\|_{2}^{2}+\alpha}+\frac{\|{\bm{\mathbf{{h}}}}_{i}\|_{2}^{2}}{\|{\bm{\mathbf{{h}}}}_{i}\|_{2}^{2}+\alpha}{\bm{\mathbf{{w}}}}_{i}\Big{)}_{+},$ (14)
${\bm{\mathbf{{h}}}}_{i}\leftarrow\Big{(}\frac{({\bm{\mathbf{{X}}}}-{\bm{\mathbf{{H}}}}{\bm{\mathbf{{W}}}}^{\sf T}+\alpha{\bm{\mathbf{{I}}}}){\bm{\mathbf{{w}}}}_{i}}{\|{\bm{\mathbf{{w}}}}_{i}\|_{2}^{2}+\alpha}+\frac{\|{\bm{\mathbf{{w}}}}_{i}\|_{2}^{2}}{\|{\bm{\mathbf{{w}}}}_{i}\|_{2}^{2}+\alpha}{\bm{\mathbf{{h}}}}_{i}\Big{)}_{+}$ (15)
@@
which are mathematically equivalent to those in Equation 13. The derivation can be found in Appendix B. This formulation allows for the updates of all the columns of ${\bm{\mathbf{{W}}}}$ then all the columns of ${\bm{\mathbf{{H}}}}$ or vice versa. Choosing to update all ${\bm{\mathbf{{w}}}}_{i}$’s followed by all ${\bm{\mathbf{{h}}}}_{i}$’s allows for the products ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}}$ and ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$ to be computed and reused through a single sweep of updates over all columns of ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ resulting in better computational efficiency in practice. To illustrate this, consider the update for ${\bm{\mathbf{{w}}}}_{i}$. The bulk of the computation is needed for computing $({\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}+\alpha{\bm{\mathbf{{I}}}}){\bm{\mathbf{{h}}}}_{i}$ which is the $i$th column of $({\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}+\alpha{\bm{\mathbf{{H}}}})$, where the product ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$ will not change as each ${\bm{\mathbf{{w}}}}_{i}$ is updated. The same applies for the columns of ${\bm{\mathbf{{H}}}}$. Overall our proposed updates shown in Equation 14 are more memory efficient and more computationally efficient by a factor of 2.

### v2

@@ Abstract
Symmetric Nonnegative Matrix Factorization (SymNMF) is a technique in data analysis and machine learning that approximates a symmetric matrix with a product of a nonnegative, low-rank matrix and its transpose. To design faster and more scalable algorithms for SymNMF we develop two randomized algorithms for its computation. The first algorithm uses randomized matrix sketching to compute an initial low-rank approximation to the input matrix and proceeds to rapidly compute a SymNMF of the approximation. The second algorithm uses randomized leverage score sampling to approximately solve constrained least squares problems. Many successful methods for SymNMF rely on (approximately) solving sequences of constrained least squares problems. We prove theoretically that leverage score sampling can approximately solve nonnegative least squares problems to a chosen accuracy with high probability. Additionally, we prove sampling complexity results for previously proposed hybrid sampling techniques which deterministically include high leverage score rows. This hybrid scheme is crucial for obtaining speeds ups in practice. Finally we demonstrate that both methods work well in practice by applying them to graph clustering tasks on large real world data sets. These experiments show that our methods approximately maintain solution quality and achieve significant speed ups for both large dense and large sparse problems.
keywords:
Nonnegative Matrix Factorization, Symmetric Nonnegative Matrix Factorization, Leverage Scores, matrix sketching, randomized numerical linear algebra (RandNLA)
{AMS}
05C50, 65F20, 65F55, 65F50, 90C20
@@ 1 Introduction
We propose the first randomized algorithms for Symmetric Nonnegative Matrix Factorization (SymNMF). Nonnegative Matrix Factorization
@@
(NMF) is an important method in data analysis with applications to data visualization, text mining, feature learning, information fusion and more [38, 35, 57, 31, 16]. SymNMF is a variant of NMF where the input matrix is symmetric and the output low-rank approximation is also constrained to be symmetric [35, 61]. Applications of SymNMF include (hyper)graph clustering, image segmentation, and information fusion [55, 15, 27, 7, 8]. Several randomized algorithms for nonsymmetric NMF have been previously proposed and shown to be effective for dense and small sparse problems [54, 51, 19], but as far as we are aware there is no prior work on randomized algorithms for SymNMF. Along the way we also prove two interesting results generalizing existing guarantees for leverage score sampling for overdetermined least squares problems to all convex overdetermined least squares problems (e.g. nonnegative least squares) and prove theoretical guarantees for hybrid leverage score sampling schemes applied to such problems. Our contributions include:
  • a randomized algorithm for SymNMF we call “Low-rank Approximated Input SymNMF” (LAI-SymNMF),
  • a randomized algorithm based on leverage score sampling for least squares problems we call LvS-SymNMF,
  • novel theoretical analysis of leverage score sampling for the Nonnegative Least Squares problem and more generally convex least squares problems,
  • theoretical analysis of a hybrid leverage score sampling scheme for convex least squares problem,
  • and experiments on two large, real world clustering tasks.
@@
The rest of the paper is organized as follows. Section 2, which discusses background material including non-randomized SymNMF algorithms, reviews existing randomized NMF methods and other related work such as randomized methods for other low-rank matrix and tensor decompositions. Section 3 introduces our first proposed algorithm LAI-SymNMF. LAI-SymNMF uses randomized methods to rapidly compute an initial, unconstrained low-rank approximation before proceeding to compute an NMF of this LAI. Section 4 presents a row-sampling algorithm called LvS-SymNMF. This method solves a sequence of nonnegative least squares (NLS) problems using a technique called leverage score sampling to accelerate the solver. Additionally, we use a hybrid approach from [36] that involves both deterministic and randomized sampling based on the leverage scores. Our novel theoretical analysis of this hybrid strategy gives the sample complexity needed to achieve an accuracy guarantee relative to the NLS residual, with high probability, and we empirically show its advantage over purely randomized sampling. Section 5 presents experimental results for the proposed algorithms on two real world data sets. Each data set is represented as a graph and the SymNMF output is used to cluster the vertices. Both methods achieve significant speed ups over deterministic methods ranging from $5$ to 7.5×7.5$\times$7.5 × and are able to maintain accuracy in terms of normalized residual norms and cluster quality.
2 Preliminaries
@@
We begin by briefly discussing the NMF and SymNMF problems followed by an introduction to various methods for computing NMF’s and SymNMF’s. Then we discuss the “Randomized Range Finder” and leverage score sampling for least squares problems. The section also contains a statement of Theorem 2.1 which provides novel bounds for Leverage Score sampling applied to Nonnegative Least Squares Problems. Various notation is given in Table 1.
@@
${\bm{\mathbf{{I}}}}$ Identity Matrix ${\bm{\mathbf{{a}}}}_{i}$ $i$th column vector of ${\bm{\mathbf{{A}}}}$
@@
${\bm{\mathbf{{A}}}}^{\sf T}$ Transposition $\mathbb{R}_{+}$ Nonnegative Real Numbers
@@
$\sigma_{i}({\bm{\mathbf{{A}}}})$ $i$th largest singular value of ${\bm{\mathbf{{A}}}}$ $\sigma_{max}({\bm{\mathbf{{A}}}}),\sigma_{min}({\bm{\mathbf{{A}}}})$ largest and smallest singular value
@@
(1) $\min_{\{{\bm{\mathbf{{W}}}},{\bm{\mathbf{{H}}}}\}\geq 0}\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}$
where ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{m\times n}$, ${\bm{\mathbf{{W}}}}\in\mathbb{R}_{+}^{m\times k}$, and ${\bm{\mathbf{{H}}}}\in\mathbb{R}_{+}^{n\times k}$. The notation $\mathbb{R}_{+}$ denotes the nonnegative orthant and $k$ is the desired reduced rank, given as an input, and usually $k\ll\min(m,n)$. While ${\bm{\mathbf{{X}}}}$ is often also nonnegative, it is not strictly required. Since NMF is a non-convex optimization problem, optimization algorithms often find a local minimum at best. The works [23, 32] give a comprehensive discussion of NMF.
When ${\bm{\mathbf{{X}}}}$ is symmetric, ${\bm{\mathbf{{X}}}}={\bm{\mathbf{{X}}}}^{\sf T}$, it is often desired the two low-rank factors in Equation 1 be the same [35, 52, 61, 34]. This problem is called Symmetric NMF (SymNMF) and its objective function is expressed as
(2) $\min_{{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{H}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}.$
@@ 2.1 Algorithms for SymNMF
Many algorithms for solving SymNMF shown in Equation 2 have been proposed. Broadly these methods can be put into two categories: Alternating Updating (AU) methods and all-at-once optimization methods. AU methods alternate between updating a subset of the variables while holding others fixed, eventually iterating through the entire subset of variables. All-at-once methods update all of the variables simultaneously.
AU methods for SymNMF include symmetrically regularized Alternating Nonnegative Least Squares (ANLS) [35], symmetrically regularized Hierarchical Least Squares (HALS) [61], Cyclic Coordinate Descent (CCD) [52], and Progressive Hierarchical Alternating Least Squares (PHALS) [28]. All-at-once methods include Projected Gradient Descent (PGD), Projected Newton-like update [35], and Projected Gauss-Newton with Conjugate Gradients (PGNCG) [22]. We focus on methods based on regularized ANLS and the PGNCG algorithm. This is because the ANLS method is generally superior to both PGD and the Newton-like method [35], and the CCD method tends to be unsuitable for large data sets as it sequentially iterates over the elements of ${\bm{\mathbf{{H}}}}$. We now present the methods based on regularized ANLS and the PGNCG algorithm.
@@ 2.1.1 Symmetrically Regularized Alternating Nonnegative Least
(3) $\min_{\{{\bm{\mathbf{{W}}}},{\bm{\mathbf{{H}}}}\}\geq 0}\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}^{2}+\alpha\|{\bm{\mathbf{{W}}}}-{\bm{\mathbf{{H}}}}\|_{F}^{2},$
where ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ are forced to be close to each other in Frobenius norm using a large value of $\alpha$, thus guiding the iterates towards a symmetric approximation [35]. The authors of [61] show that under mild assumptions the critical points of Equation 2 and Equation 3 are the same and argue that this makes Equation 3 an appropriate surrogate for Equation 2. Equation 3 can be iteratively approximated by using the ANLS method for updating ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ as in the following two equations :
(4) $\min_{{\bm{\mathbf{{W}}}}\geq 0}\Big{\|}\begin{bmatrix}{\bm{\mathbf{{H}}}}\\ \sqrt{\alpha}{\bm{\mathbf{{I}}}}\end{bmatrix}{\bm{\mathbf{{W}}}}^{\sf T}-\begin{bmatrix}{\bm{\mathbf{{X}}}}\\ \sqrt{\alpha}{\bm{\mathbf{{H}}}}^{\sf T}\end{bmatrix}\Big{\|}_{F}\ \text{and}\ \min_{{\bm{\mathbf{{H}}}}\geq 0}\Big{\|}\begin{bmatrix}{\bm{\mathbf{{W}}}}\\ \sqrt{\alpha}{\bm{\mathbf{{I}}}}\end{bmatrix}{\bm{\mathbf{{H}}}}^{\sf T}-\begin{bmatrix}{\bm{\mathbf{{X}}}}\\ \sqrt{\alpha}{\bm{\mathbf{{W}}}}^{\sf T}\end{bmatrix}\Big{\|}_{F}.$
@@
The ANLS method is generally superior to PGD and the Newton-like method [35]. This is mainly because PGD suffers from slow convergence and the Newton-like method is expensive for even moderately sized problems due to the need for approximate Hessian inversion. To solve the ANLS formulation we use the Block Principle Pivoting (BPP) solver from [33].
@@ 2.1.2 Hierarchical Alternating Least Squares for SymNMF
The HALS method for NMF was developed in [24, 9]. A method that uses the HALS framework for optimizing Equation 3 was proposed in [61]. The update rules are given by the following two equations
(5) ${\bm{\mathbf{{w}}}}_{i}=\max\Big{(}\frac{({\bm{\mathbf{{R}}}}_{i}+\alpha{\bm{\mathbf{{I}}}}){\bm{\mathbf{{h}}}}_{i}}{\|{\bm{\mathbf{{h}}}}_{i}\|_{2}^{2}+\alpha},0\Big{)}\ \text{and}\ {\bm{\mathbf{{h}}}}_{i}=\max\Big{(}\frac{({\bm{\mathbf{{R}}}}_{i}+\alpha{\bm{\mathbf{{I}}}}){\bm{\mathbf{{w}}}}_{i}}{\|{\bm{\mathbf{{w}}}}_{i}\|_{2}^{2}+\alpha},0\Big{)}.$
@@
Here ${\bm{\mathbf{{W}}}}=\big{[}{\bm{\mathbf{{w}}}}_{1},\cdots,{\bm{\mathbf{{w}}}}_{k}\big{]}$, ${\bm{\mathbf{{H}}}}=\big{[}{\bm{\mathbf{{h}}}}_{1},\cdots,{\bm{\mathbf{{h}}}}_{k}\big{]}$ are the columns of ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ and ${\bm{\mathbf{{R}}}}_{i}={\bm{\mathbf{{X}}}}-\sum_{j\neq i,j=1}^{k}{\bm{\mathbf{{w}}}}_{j}{\bm{\mathbf{{h}}}}_{j}^{\sf T}.$ Following these rules, the columns of ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ are updated as pairs in sequence as ${\bm{\mathbf{{w}}}}_{1},{\bm{\mathbf{{h}}}}_{1},\cdots{\bm{\mathbf{{w}}}}_{i},{\bm{\mathbf{{h}}}}_{i},\cdots{\bm{\mathbf{{w}}}}_{k},{\bm{\mathbf{{h}}}}_{k}$. This method is less efficient as it relies on the matrix ${\bm{\mathbf{{R}}}}_{i}$, as in Equation 5. To make the updates efficient we modify the update rule following algorithms from [24] for more efficient HALS updates to
(6) ${\bm{\mathbf{{w}}}}_{i}\leftarrow\Big{(}\frac{({\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}+\alpha{\bm{\mathbf{{I}}}}){\bm{\mathbf{{h}}}}_{i}}{\|{\bm{\mathbf{{h}}}}_{i}\|_{2}^{2}+\alpha}+\frac{\|{\bm{\mathbf{{h}}}}_{i}\|_{2}^{2}}{\|{\bm{\mathbf{{h}}}}_{i}\|_{2}^{2}+\alpha}{\bm{\mathbf{{w}}}}_{i}\Big{)}_{+},$
(7) ${\bm{\mathbf{{h}}}}_{i}\leftarrow\Big{(}\frac{({\bm{\mathbf{{X}}}}-{\bm{\mathbf{{H}}}}{\bm{\mathbf{{W}}}}^{\sf T}+\alpha{\bm{\mathbf{{I}}}}){\bm{\mathbf{{w}}}}_{i}}{\|{\bm{\mathbf{{w}}}}_{i}\|_{2}^{2}+\alpha}+\frac{\|{\bm{\mathbf{{w}}}}_{i}\|_{2}^{2}}{\|{\bm{\mathbf{{w}}}}_{i}\|_{2}^{2}+\alpha}{\bm{\mathbf{{h}}}}_{i}\Big{)}_{+}$
@@
which are mathematically equivalent to those in Equation 5. These rules can be seen as combining the update order for HALS from [9] and regularization, such as used in [10]. A derivation can be found in the Appendix A. This formulation allows for the updates of all the columns of ${\bm{\mathbf{{W}}}}$ then all the columns of ${\bm{\mathbf{{H}}}}$ or vice versa. As discussed in [24], choosing to update all ${\bm{\mathbf{{w}}}}_{i}$’s followed by all ${\bm{\mathbf{{h}}}}_{i}$’s allows for the products ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}}$ and ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$ to be computed and reused through a single sweep of updates over all columns of ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ resulting in better computational efficiency in practice. To illustrate this, consider the update for ${\bm{\mathbf{{w}}}}_{i}$. The bulk of the computation is needed for computing $({\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}+\alpha{\bm{\mathbf{{I}}}}){\bm{\mathbf{{h}}}}_{i}$ which is the $i$th column of $({\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}+\alpha{\bm{\mathbf{{H}}}})$, where the product ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$ will not change as each ${\bm{\mathbf{{w}}}}_{i}$ is updated. The same applies for the columns of ${\bm{\mathbf{{H}}}}$. Overall our proposed updates shown in Equation 6 are more memory efficient and more computationally efficient by a factor of 2.